OpenStudy (anonymous):

Domain and range of -9/(sqrt z+1)

mathslover (mathslover):

Is it : \(\cfrac{-9}{\sqrt{z+1}} \) or \(\cfrac{-9}{\sqrt{z} + 1}\) ?

OpenStudy (mathmale):

A solution might be slightly easier to visualize if you'd type your expression in this way:\[\frac{ -9 }{ \sqrt{z+1} }\] What restrictions on z do you see?

OpenStudy (mathmale):

mathslover poses a valid question. Which of the two expressions did you mean to type in?

mathslover (mathslover):

@OscarIng - If you want us to help you appropriately, please let us know what exactly the question is.